# Lindblad-like quantum tomography for non-Markovian quantum dynamical maps

### v1

@@ III Lindblad-like quantum tomography for non-Markovian dephasing
Here, we have rewritten this measure in terms of the error probability of the phase-flip channel (23), which changes infinitesimally with $\dot{p}(t)=\gamma(t)\mathord{\rm e}^{-\Gamma(t)}$. Hence, the non-Markovianity condition translates into a dynamical situation in which phase-flip errors do not increase monotonically during the whole evolution. When the dephasing rate attains negative values, the phase-flip error probability can decrease, such that the qubit momentarily recoheres (see the orange lines in Fig. 1). In this simple case of pure dephasing noise, we see that both measures of non-Markovianity in Eqs. (24) and  (25) depend on the rate $\gamma(t)$ attaining negative values, and both are equal to 0 if $\gamma(t)$ is always positive.
Once these additional properties of the dephasing quantum dynamical map have been discussed, we can move back to the estimation $\ell$QT protocol (11), and how it can be simplified even further. As noted above, in contrast to the informationally-complete set of initial states an measurements that must be considered for the general cost function of $\ell$QT (9), we can work with a smaller number of configurations by noting that all information of the dephasing map can be extracted by preparing a single initial state $\rho_{0}=\outerproduct{+}{+}$ |, and measuring in a single basis $M_{x,\pm}$ (see Fig. 1). As advanced in the introduction, we would like to know how many snapshots $|\mathbb{I}_{t}|$ are required, at which the system is measured after evolving for $\{t_{i}:\,i\in\mathbb{I}_{t}\}$ and, moreover, which are the optimal times of those snapshots in terms of the specific details of the noise. We take here two different routes. On the one hand, we can follow a frequentist approach similarly to the one in Markovian Lindblad quantum tomography [87, 89, 90], but now taking into account the time-dependence of the decay rates, which will require an a priori selection of the evolution times at which the system is probed. Alternatively, we also consider a Bayesian approach that exploits a physically-motivated prior knowledge about the noise parameters, which can be represented by a certain probability distribution. After performing measurements on the system at certain instants of time, we update this probability distribution with the new acquired information, a process that is repeated until reaching a target accuracy for the parameter estimation. The Bayesian approach has the advantage of choosing, at each step, the most convenient subsequent time at which to measure by maximizing the information one would gain. In principle, this can lead to a reduction in the number of measurements required to reach a certain accuracy with respect to those required by the frequentist approach. In practice, however, the non-Markovianity of the quantum evolution can modify this argument, as our probabilistic account of the model parameters can be affected by the actual time correlations of the underlying random process. We present below a detailed comparison of the two approaches, determining the regimes in which each of them is better than the other.
III.1 Frequentist approach to non-Markovian inference
In this section, we focus on the frequentist approach, which builds on the relative frequencies of observed outcomes. The $\ell$QT cost function (9) can be rewritten as follows
@@ III.1 Frequentist approach to non-Markovian inference
which is greatly simplified with respect to the general case in Eq. (5), as we only have a single initial state and a single measurement basis. Since we are monitoring the coherence of the qubit, this cost function corresponds to a Ramsey-type estimator, where $\tilde{f}_{i}(0\,|\bm{\theta}_{\star})=N_{i,0}/N_{i}$ $(\tilde{f}_{i}(1\,|\bm{\theta}_{\star})=N_{i,1}/N_{i})$ is the ratio of the number of outcomes observed $N_{i,0}$ ($N_{i,1}=N_{i}-N_{i,0}$) to the total of $N_{i}$ shots collected at the instant of time $t_{i}$, when measuring the system with the POVM element $M_{x,0}$ ($M_{x,1}$). Our notation remarks that these relative frequencies carry information about the real noise parameters $\bm{\theta}_{\!\star}$ we aim at estimating. In addition, the estimator depends on $p_{i}(m_{x}|\bm{\theta})$ (17)-(18), which stand for the probabilities obtained by solving the time-local dephasing master equation (12), where we make explicit the dependence on the parametrized noise.
The minimization (11) is simplified in this case, as one can dispense with the time-dependent constraints. As a consequence, the frequentist approach can be recast as a statistical problem of parameter point estimation [96], namely
$\bm{\hat{\theta}}_{\rm F}=\texttt{argmin}_{\bm{\theta}}\{\mathsf{C}^{\rm pd}_{\mathrm{TL}}(\bm{\theta}):\,\bm{\theta}\in\Theta=\mathbb{R}^{n}\}.$ (27)
Instead of the general $\ell$QT learning over $d^{2}(d^{2}-1)$ parameters, which increase exponentially $d=2^{n_{q}}$ with the number of qubits $n_{q}$ and can arbitrarily change in time, our procedure revolves around the estimation of $n$ noise parameters, which are independent of the system size and the evolution times. On the other hand, the imprecision of our estimates will indeed depend on our choice of the evolution times, forcing us to go beyond the LQT single-time estimator [90], and actually measure at optimal times for which the estimation imprecision can be minimized. Let us note that the conditions to minimise this cost function are the same as those that minimize the Kullback-Leibler divergence [103, 104], which is the following relative entropy
between the experimental $p_{1,k}\in\{\tilde{f}_{i}(+\,|\bm{\theta}_{\star}),(1-\tilde{f}_{i}(+\,|\bm{\theta}_{\star}))\}$ and parametrized theoretical $p_{2,k}\in\{p^{\rm TL}_{i}(+\,|\bm{\theta}),p^{\rm TL}_{i}(-\,|\bm{\theta})\}$ probability distributions, provided one considers variations with respect to the estimation parameters $\bm{\theta}$.
We note that the expectation values are taken with respect to the probability distributions for the measurements (3) which, implicitly, also have the stochastic average over the random dephasing noise in the semi-classical model, or a partial trace over the environment/bath in the quantum-mechanical one. The nice property of the maximum-likelihood estimator is that it is asymptotically unbiased, such that $\bm{\mathsf{B}}_{\bm{\theta}_{\star}}\!\!(\bm{\theta}_{\rm F})=\mathbb{E}[\bm{\hat{\theta}}_{\rm F}]-\bm{\theta}_{\star}\to\bm{0}$ for a sufficiently-large $N_{\rm shot}\gg 1$. Moreover, its asymptotic covariance matrix saturates the Cramér-Rao bound [105] relating the estimation precision to the Fisher information matrix, which quantifies the amount of information in $\mathbb{D}$ about the unknown parameters. If we momentarily assume that the measurements occur at a single instant of time $t=t_{i}$ with outcomes that are independent and identically distributed, the covariance matrix becomes $\mathrm{Cov}(\bm{\hat{\theta}}_{\rm F})\approx(N_{i}{I}_{F,i}(\bm{\theta}_{\!\star}))^{-1}$, where
$[{I}_{F,i}(\bm{\theta}_{\!\star})]_{n_{1},n_{2}}=\mathbb{E}\left[\left.\frac{\partial{\rm log}\big{(}p_{i}^{\rm TL}(m_{x}|\bm{\theta})\big{)}}{\partial{\theta_{n_{1}}}}\frac{\partial{\rm log}\big{(}p_{i}^{\rm TL}(m_{x}|\bm{\theta})\big{)}}{\!\partial\theta_{n_{2}}}\!\right|_{\bm{\theta}_{\star}}\right].$ (30)
In this work, we deal with the more general case in which we measure at several times, therefore the random variables are not identically distributed, and the total number of shots need not be the same for different times. In this case, we must take a linear combination of the Fisher information matrices of each measurement time weighted by the proportion of measurements taken at each time  [106]. We thus obtain the asymptotic covariance matrix
$\mathrm{Cov}(\bm{\hat{\theta}}_{\rm F})\approx\Sigma_{\bm{\hat{\theta}}}\equiv\left[N_{\rm shot}\sum_{i}\frac{N_{i}}{N_{\rm shot}}{I}_{F,i}(\bm{\theta}_{\!\star})\right]^{-1},$ (31)
such that, the more the Ramsey estimator varies under changes of the noise parameters, the bigger the amplification of the noise parameter is and, thus, the smaller the imprecision one can achieve. As we can see, the imprecision of the estimate will scale with $1/\sqrt{N_{\rm shot}}$, such that the Ramsey estimator is asymptotically consistent in the $N_{\rm shot}\gg 1$ asymptotic limit [96]. The asymptotic statistics of the maximum-likelihood estimator is further explained in App. D.
We also note that, in this limit, the observed relative frequencies will be normally distributed, such that one can consider minimizing the weighted least-squares cost function
$\mathsf{C}_{\mathrm{TL}}^{\rm pd}(\bm{\theta})\approx\sum_{i}\frac{\big{(}\tilde{f}_{i}(0\,|\bm{\theta}_{\star})-p_{i}^{\rm TL}(0|\bm{\theta})\big{)}^{2}}{\tilde{\sigma}^{2}_{f_{i}}},$ (32)
where $\tilde{\sigma}^{2}_{f_{i}}$ is the variance of the measured samples. For large $N_{i}$ the expected variance of the measurements at time $t_{i}$ is
$\sigma^{2}_{f_{i}}=\frac{p_{i}^{\rm TL}(0\,|\bm{\theta})\big{(}1-p_{i}^{\rm TL}(0\,|\bm{\theta})\big{)}}{N_{i}},$ (33)
since we are sampling from a binomial distribution. This approximation allows us to use a simpler weighted least-squares algorithm, such as the trust-region reflective algorithm implemented in SciPy, where we can optionally set some bounds for the parameters to be estimated [107].
Once the properties of the Ramsey frequentist estimate $\bm{\hat{\theta}}_{\rm F}$ have been discussed, we can search for the optimal measurement times that would lead to a Ramsey estimator with the lowest possible imprecision for a given finite $N_{\rm shot}$. Depending on which parameter we are interested in, we may be interested in minimizing a particular component of the covariance matrix $[\Sigma_{\bm{\hat{\theta}}}]_{n_{1},n_{1}}$ (31) or, alternatively, minimize its determinant as a whole. In the asymptotic limit in which $\bm{\hat{\theta}}_{\rm F}$ follows a multivariate normal distribution, $\det\Sigma_{\bm{\hat{\theta}}}$ is proportional to the volume enclosed by the covariance elliptical region, so it is a good measure of the dispersion of the distribution, and a good way to quantify the accuracy of the estimation. Sometimes, we will also use $\det\Sigma_{\bm{\hat{\theta}}}^{1/2n}$ instead, which can be more easily compared to the individual standard deviations $[\Sigma_{\bm{\hat{\theta}}}]_{n_{1},n_{1}}^{1/2}$, as both quantities scale with $1/\sqrt{N_{\rm shot}}$. Minimizing $\det\Sigma_{\bm{\hat{\theta}}}$ has also the advantage that the optimal measurement times obtained are independent of the parameters we want to determine, assuming the different parametrizations have the same number of parameters and that a coordinate transformation exists between parametrizations. In this case, the determinants of the covariance matrices are related by $\det\Sigma_{\bm{\hat{\theta}}}=\det\Sigma(\bm{\hat{\theta}}^{\prime}_{\rm F})\det J^{2}$, with $J$ the Jacobian of the coordinate transformation between both parametrizations, which does not depend on time and therefore will have no influence in the minimization. Before presenting these results, we discuss an approach based on Bayesian inference [96].
III.2 Bayesian approach to non-Markovian inference
Rather than considering the relative frequencies as approximations of the underlying probability distribution with a certain fixed value of $\bm{\theta}_{\star}$, the idea of Bayesian inference is to quantify statistically our knowledge about the noise parameters, and how this knowledge gets updated as we collect more information via measurements. Hence, the noise parameters become continuous stochastic variables themselves $\bm{\theta}_{\star}\mapsto\bm{\tilde{\theta}}$ that take values $\bm{\theta}\in\Theta$ according to a prior pdf $\pi_{0}(\bm{\theta})$. This probability distribution quantifies our uncertainty about the noise parameters before making any measurement $\mathbb{D}_{0}=\emptyset$. At each $\ell>0$ Bayesian step, we measure the system enlarging the data set sequentially $\mathbb{D}_{\ell-1}\mapsto\mathbb{D}_{\ell}=\mathbb{D}_{\ell-1}\cup\delta\mathbb{D}_{\ell}$, where $\delta\mathbb{D}_{\ell}\subset\mathbb{D}=\{N_{i,+,x,m_{x}}\}$ contains a number of measurement outcomes $|\delta N_{\ell}|$ that is a fraction of the total $N_{\rm shot}$. These outcomes will be labelled as $\delta\mathbb{D}_{\ell}=\{N_{i_{\ell},+,x,m_{x_{\ell}}}\}$. The measurements in this data set are again binary Bernoulli trials, and can be described by a joint multinomial distribution $p_{\tilde{N}}(\delta\mathbb{D}_{\ell})$ defined in analogy to Eq. (4), but only extended to the configurations measured in the particular Bayesian step. The prior $(\ell-1)$-th probability distribution is then updated by using Bayes’ rule based on the parametrized probability distributions $p^{\rm TL}_{\ell}(\delta{\mathbb{D}}_{\ell}|\bm{\theta})$ being understood as probabilities conditioned on our statistical knowledge of the noise parameters
[…]
$\mathsf{R}_{\rm TL}(\hat{\tau})=\!\sum_{i_{\ell},m_{x_{\ell}}\in\delta\mathbb{D}_{\ell}}\int_{\Theta}\!{\rm d}^{n}\theta\,\pi_{\ell}(\bm{\theta})p^{\rm TL}_{\ell}\big{(}\delta\mathbb{D}_{\ell}|\bm{\theta}\big{)}(\hat{\tau}-\bm{\theta})^{2}.$ (36)
In addition to the expectation value (35), since we have the updated probability distributions, we can quantify how our uncertainty about the noise parameters changes via the associated covariance matrix $[\mathrm{Cov}(\bm{\hat{\theta}}_{{\rm B}})]_{n_{1},n_{2}}$ or any other statistics, regardless of the size of the Bayesian data set $\mathbb{D}_{\ell}$. This differs from our previous arguments for covariance of the maximum likelihood estimate $\bm{\hat{\theta}}_{\rm F}$ (31), which require working in the asymptotic regime $N_{\rm shot}\gg 1$. In experimental situations in which this regime cannot be reached, we note that one could use Monte Carlo sampling techniques to estimate precision of $\bm{\hat{\theta}}_{\rm F}$ [108, 109, 77], although these deal with the estimator based on the full likelihood function in Eq. (4).
Another crucial difference of the Bayesian approach is that, instead of choosing a predefined set of evolution times $\{t_{i}:\,\in\mathbb{I}_{t}\}\subset T$, either distributed uniformly, randomly, or at precise instants to minimise the covariance, we can find the optimal time at which we should measure to maximize the information gain at each Bayesian step. For each update, we thus solve for
$t_{\ell}=\texttt{argmax}_{t}\big{\{}\mathbb{E}\left[\mathsf{D}_{\mathrm{KL}}\big{(}\pi_{\ell}(\bm{\theta}|\mathbb{D}_{\ell})\,\big{|}\!\big{|}\,\pi_{\ell}(\bm{\theta})\big{)}\right]\big{\}},$ (37)
where $\pi_{\ell}(\bm{\theta}|\mathbb{D}_{\ell})$ is the posterior probability (34) corresponding to the measurement results one would obtain by measuring at a time $t$ and enlarging the data set as $\mathbb{D}_{\ell^{\prime}}=\mathbb{D}_{\ell}\cup\delta\mathbb{D}_{\ell^{\prime}}$. In the expression above, we are making use of the Kullback-Leibler divergence (28) between the posterior and the prior, searching for a time that maximizes the relative entropy between the prior and any of the possible posteriors, such that one gains the maximum amount of information at each Bayesian step. Therefore, not only the data set is enlarged sequentially $\mathbb{D}_{\ell}\supset\mathbb{D}_{\ell-1}$, but also the specific times are chosen adaptively. In light of the fixed set of measurement times in the frequentist approach $\{t_{i}:\,i\in\mathbb{I}_{t}\}$, we note that the total set of updated times after $N_{\rm B}$ Bayesian steps $\{t_{\ell},\ell\in\{0,\cdots,N_{\rm B}\}\}$ can be very different, and that is the reason why we use a different notation. Computing Eq. (37) can be quite time-consuming, especially when dealing with a large number of parameters. In practice, long computation times may lead to a reduction in the frequency of experimental shots, which is undesirable. To avoid this, we can take tenths or hundreds of shots at each step before computing again the optimal measurement time. This will not change the results significantly, since a Bayesian update of a single shot does not change the prior much and the optimal measurement time of next step remains very similar to the previous one.
We note that the Bayesian approach is ultimately related to the maximum-likelihood estimation in the asymptotic limit $N_{\rm shot}\gg 1$. When the variance of the prior is small, the maxima of the Kullback-Leibler divergence (28) between prior and posterior are localized at the optimal times obtained by minimizing the asymptotic covariance matrix of the maximum likelihood-estimation (31). Additionally, if we make a single Bayesian update in which we take a very big number of shots at several times, the likelihood function relating the posterior and the prior will be a multinomial pdf (4). Given the big number of shots taken in this single step, the posterior distribution will be mainly shaped by the likelihood function, which contains most of the information about the parameters. The position of the maximum of this pdf will be located at the most likely value of the parameters, and therefore it is in agreement with the maximum-likelihood estimator that takes measurements at those times. The Bayesian approach has the advantage that, at each step, we measure at the most convenient time that maximizes the expected information gain, and this can lead to an overall reduction of the number of measurements needed. Also, as noted above, the final estimate relies on a probability distribution, so that we can immediately derive confidence intervals without requiring any asymptotic limit. For the Bayesian protocol design, we have used the Python package Qinfer [110], which numerically implements the operations needed by using a sequential Monte Carlo algorithm for the updates.
III.3  Comparative performance analysis
Let us now study some dephasing dynamics in which we can apply the two approaches we have just introduced, and make a comparative study of their performance when learning parametrized dephasing maps with time-correlated noise.
III.3.1 Markovian semi-classical dephasing
@@
where $c>0$ is a so-called diffusion constant. Since this correlations only depend on the time differences, the process is wide-sense stationary. Moreover, on the basis of its Gaussian joint pdf, it can be shown that the process is indeed strictly stationary. Being Gaussian, all the information is thus contained in its two-point functions or, alternative, in its PSD (59)
@@
which has a Lorentzian shape. This Gaussian process then leads to an exact time-local master equation for the dephasing of the qubit (12) with a time-dependent decay rate
@@
In the long-time limit $t\gg\tau_{\rm c}$, one recovers a constant decay rate $\gamma(t)\approx S(0)/4=c\tau_{\rm c}^{2}/4$, which connects to our previous discussion of the effective exponential decay of the Ramsey signal $p^{\rm TL}_{i}(m_{x})\approx\big{(}1+(-1)^{m_{x}}\mathord{\rm e}^{-t_{i}/T_{2}}\big{)}/2$ and the decoherence time $T_{2}=2/c\tau_{\rm c}^{2}$. On the other hand, for shorter time scales, we see the effects of the noise memory through a time-inhomogeneous evolution of the coherences that goes beyond a Lindbladian description. It is interesting to remark that, in spite of not being a Lindbladian evolution, the time-dependent decay rate is always positive, such that the two measures of non-Markovianity in Eqs. (24)-(25) vanish exactly. The pure dephasing quantum dynamical map of a qubit subjected to OU frequency noise is thus Markovian albeit not Lindbladian.
@@
Frequentist Ramsey estimators.
We can now evaluate the $\ell$QT cost function $\mathsf{C}^{\rm pd}_{\mathrm{TL}}(\bm{\theta})$ in Eq. (26) by substituting the attenuation factor (41) in the likelihood function $p^{\rm TL}_{i}(m_{x}|\bm{\theta})=\big{(}1+(-1)^{m_{x}}\mathord{\rm e}^{-\Gamma(t_{i})}\big{)}/2$ after a certain set of evolution times $\{t_{i},i\in\mathbb{I}_{t}\}$, and the relative frequencies for the measurement outcomes $\tilde{f}_{i}(m_{x}\,|\bm{\theta}_{\star})$. For a Lindbladian dynamics, a single measuring time $t_{1}$ would suffice for the estimation [90], which can actually be solved for analytically in the present pure dephasing context, as discussed in App. B. For the OU dephasing, this is no longer the case, and we actually need at least two times $t_{1},t_{2}$. In order to assess the performance of the frequentist minimization problem (27) under shot noise, we numerically generate the relative frequencies $\tilde{f}_{1}(m_{x}\,|\bm{\theta}_{\star}),\tilde{f}_{2}(m_{x}\,|\bm{\theta}_{\star})$ at two instants of time by sampling the probability distribution with the actual OU parameters $\bm{\theta}_{\star}=(c_{\star},\tau_{\rm c{\star}})$ a number of times $N_{\rm shot}=N_{1}+N_{2}$. In the following, rather than learning $(c_{\star},\tau_{\rm c\star})$, we will focus on two noise parameters with units of time $(T_{2\star},\tau_{\rm c\star})$, where we recall that $T_{2\star}=2/c^{\phantom{2}}_{\star}\tau_{{\rm c}\star}^{2}$ is an effective decoherence time in the long-time limit. In Fig. 2, we present a contour plot of this two-time cost function $\mathsf{C}^{\rm pd}_{\mathrm{TL}}(\bm{\theta})$, which is actually convex and allows for a neat visualization of its global minimum. We also depict with a red cross the result of a gradient-descent minimization, where one can see that the estimates $\bm{\hat{\theta}}_{\rm F}=(\hat{T}_{2},\hat{\tau}_{\rm c})$ are close to the real noise parameters. The imprecision of the estimate is a result of the shot noise, which we now quantify.
Figure 2: Gradient descent for OU dephasing $\ell$QT: Contour plot of the Ramsey cost function $\mathsf{C}^{\rm pd}_{\mathrm{TL}}(\bm{\theta})$ of Eq. (26) as a function of the noise parameters $(T_{2},\tau_{\rm c})$, where we recall that $T_{2}=2/c\tau_{\rm c}^{2}$ plays the role of an effective decoherence time in the long-time limit. We choose two evolution times and, for illustration purposes, fix them at $t_{1}=\tau_{\rm c\star}$, $t_{2}=2\tau_{\rm c\star}$, fixing the OU diffusion constant such that $T_{2\star}/\tau_{\rm c\star}=1$. We distribute $N_{\rm shot}=2\times 10^{3}$ shots equally per time step. The gradient descent of this convex problem converges towards the global minimum, and is marked with a red cross that lies close to the real noise parameters $\bm{\theta}\mapsto\bm{\theta}_{\star}$.

### v2

@@ Results
Here, we have rewritten this measure in terms of the error probability of the phase-flip channel of Eq. (31), which changes infinitesimally with $\dot{p}(t)=\gamma(t)\mathord{\rm e}^{-\Gamma(t)}$. Hence, the non-Markovianity condition translates into a dynamical situation in which phase-flip errors do not increase monotonically during the whole evolution. When the dephasing rate attains negative values, the phase-flip error probability can decrease, such that the qubit momentarily recoheres (see the orange lines in Fig. 1). In this simple case of pure dephasing noise, we see that both measures of non-Markovianity in Eqs. (32) and  (33) depend on the rate $\gamma(t)$ attaining negative values, and both are equal to 0 if $\gamma(t)$ is always positive. We define and detect non-Markovianity in this case using Eqs. (32) and (33). However, it is important to note that these measures provide sufficient, but not necessary, conditions for detecting all possible forms of non-Markovian behavior. For instance, certain non-Markovian dynamics may not yield an increase in trace distance.
Once these additional properties of the dephasing quantum dynamical map have been discussed, we can move back to the estimation $\ell$QT protocol of Eq. (7), and how it can be simplified even further. As noted above, in contrast to the informationally-complete set of initial states an measurements that must be considered for the general cost function of $\ell$QT in Eq. (5), we can work with a smaller number of configurations by noting that all information of the dephasing map can be extracted by preparing a single initial state $\rho_{0}=\outerproduct{+}{+}$ |, and measuring in a single basis $M_{x,\pm}$ (see Fig. 1). Indeed, this combination of initial state and measurement basis yields the expected value $p^{\rm TL}_{+,i,x}(m_{x})=\frac{1}{2}(1+(-1)^{m_{x}}\mathord{\rm e}^{-\Gamma(t_{i})})$. Similarly, if we chose $\rho_{0}=\outerproduct{+\mathrm{i}}{+\mathrm{i}}$ | and $M_{y,\pm}$, we would obtain $p^{\rm TL}_{+\mathrm{i},i,y}(m_{y})=\frac{1}{2}(1+(-1)^{m_{y}}\mathord{\rm e}^{-\Gamma(t_{i})})$, which yields the same information. The rest of combinations of initial states and measurement bases yield constant values and do not provide any information about the decay rates.
As advanced in the introduction, we would like to know how many snapshots $|\mathbb{I}_{t}|$ are required, at which the system is measured after evolving for $\{t_{i}:\,i\in\mathbb{I}_{t}\}$ and, moreover, which are the optimal times of those snapshots in terms of the specific details of the noise. We take here two different routes: the frequentist and the Bayesian approach. In the frequentist case, for pure dephasing we have the cost function
@@
which is greatly simplified with respect to the general case in Eq. (8), as we only have a single initial state and a single measurement basis. Since we are monitoring the coherence of the qubit, this cost function corresponds to a Ramsey-type estimator, where $\tilde{f}_{i}(0\,|\bm{\theta}_{\star})=N_{i,0}/N_{i}$ $(\tilde{f}_{i}(1\,|\bm{\theta}_{\star})=N_{i,1}/N_{i})$ is the ratio of the number of outcomes observed $N_{i,0}$ ($N_{i,1}=N_{i}-N_{i,0}$) to the total of $N_{i}$ shots collected at the instant of time $t_{i}$, when measuring the system with the POVM element $M_{x,0}$ ($M_{x,1}$). Our notation remarks that these relative frequencies carry information about the real noise parameters $\bm{\theta}_{\!\star}$ we aim at estimating. In addition, the estimator depends on $p_{i}(m_{x}|\bm{\theta})$ shown in Eqs. (25)-(26), which stand for the probabilities obtained by solving the time-local dephasing master equation of Eq. (20), where we make explicit the dependence on the parametrized noise. Minimizing $\det\Sigma_{\bm{\hat{\theta}}}$ we can determine the optimal measurement times of the estimator. For the Bayesian approach, at each $\ell>0$ Bayesian step, we measure the system enlarging the data set sequentially $\mathbb{D}_{\ell-1}\mapsto\mathbb{D}_{\ell}=\mathbb{D}_{\ell-1}\cup\delta\mathbb{D}_{\ell}$, where $\delta\mathbb{D}_{\ell}\subset\mathbb{D}=\{N_{i,+,x,m_{x}}\}$ contains a number of measurement outcomes $|\delta N_{\ell}|$ that is a fraction of the total $N_{\rm shot}$. These outcomes will be labelled as $\delta\mathbb{D}_{\ell}=\{N_{i_{\ell},+,x,m_{x_{\ell}}}\}$. At each step, we maximize the information gain of Eq. (19) to determine the optimal measurement time. As shown in Fig. 1 the scheme for $\ell$QT in the case of pure dephasing can be divided into the following steps:
a) Quantum experiment:
  1. 1.
    Initialize the reference state $\rho_{0}=\outerproduct{0}{0}$ | (see Fig. 1).
  2. 2.
    Prepare $\rho_{0,0}=\outerproduct{+}{+}$ | by a resonant $\pi/2$ pulse.
  3. 3.
    Let the qubit evolve under the pure dephasing for different evolution times $t_{i}$.
  4. 4.
    Apply a second resonant $\pi/2$ pulse, rotating the qubit to the Pauli-$X$ measurement basis.
  5. 5.
    Collect the outcomes $m_{x}\in\{0,1\}$ and compute the frequencies $\{f_{+,i,m_{x}}\}$.
b) Classical processing and frequentist approach:
  1. 1.
    Solve the time-local master equation with a parametrized power spectral density $S_{\bm{\theta}}(\omega)$ and a filter function, yielding the predicted probabilities $\{p_{+,i,m_{x}}(\bm{\theta})\}$.
  2. 2.
    From the quantum experiment we have a set of frequencies $\{f_{+,i,m_{x}}\}$ at different times $t_{i}$. The predicted probabilities and measured relative frequencies are used in a log-likelihood cost function, which is minimized to estimate noise parameters $\bm{\theta}$.
  3. 3.
    By calculating the covariance matrix, the precision of the estimation is determined.
c) Classical processing and Bayesian approach:
  1. 1.
    Solve the time-local master equation with a parametrized power spectral density $S_{\bm{\theta}}(\omega)$ and a filter function, yielding the predicted probabilities $\{p_{+,i,m_{x}}(\bm{\theta})\}$.
  2. 2.
    We have some prior knowledge of the parameters $\bm{\theta}$, which is represented by a prior probability distribution.
  3. 3.
    At each step, evolution time $t_{i}$ is chosen to maximize the information gain of the parameters. A new quantum experiment is performed with this new evolution time.
  4. 4.
    After each measurement or set of measurements the relative frequencies are used to update the prior distribution of the parameters.
We present below a detailed comparison of the two approaches, frequentist and Bayesian, determining the regimes in which each of them is better than the other. For the Bayesian protocol design, we have used the Python package Qinfer [119], which numerically implements the operations needed by using a sequential Monte Carlo algorithm for the updates.
Let us study some dephasing dynamics in which we can apply the two approaches we have just introduced, and make a comparative study of their performance when learning parametrized dephasing maps with time-correlated noise. We study two different dephasing models: a Markovian semi-classical dephasing and a non-Markovian quantum dephasing. These models are selected for several reasons. The semi-classical model is a well-established generalization of purely Lindbladian evolution, and in fact, contains the Lindbladian regime as a limiting case. This allows us to connect with previous results from LQT and Ramsey interferometry in the context of quantum sensing and quantum clocks. On the other hand, the non-Markovian case represents the simplest possible generalization of the Markovian scenario by introducing a shifted Lorentzian PSD, making it an ideal testbed for $\ell$QT. Furthermore, it has experimental relevance, as we can use a laser-cooled vibrational mode in trapped ions to obtain a fully-tunable implementation of this non-Markovian noise. These two dephasing models offer both classical and quantum noise characterization through the symmetry of the PSD, and they provide a versatile platform to demonstrate $\ell$QT.
Regarding the choice between frequentist and Bayesian inference, the majority of quantum technologies traditionally use frequentist approaches. However, Bayesian inference is becoming increasingly relevant, particularly in situations where the number of shots is limited and clock cycle times are large, such as in trapped-ion platforms. As we show below, Bayesian methods offer advantages in scenarios where physical priors are available and can provide better estimates in non-Markovian cases, where the dynamics are more complex.
For each one of the cases presented below, we have a PSD with some real parameters $\bm{\theta}_{\star}$ from which we can determine the time-dependent measurement probabilities $p^{\rm TL}_{i}(m_{x}|\bm{\theta_{\star}})$. From these probabilities, we numerically take the necessary samples to simulate the experiment by using the SciPy implementation of a binomial random variable [120].
Markovian semi-classical dephasing
@@
where $c>0$ is a so-called diffusion constant. Since this correlations only depend on the time differences, the process is wide-sense stationary. Moreover, on the basis of its Gaussian joint PDF, it can be shown that the process is indeed strictly stationary. Being Gaussian, all the information is thus contained in its two-point functions or, alternatively, in its PSD
@@
which has a Lorentzian shape. This Gaussian process then leads to an exact time-local master equation for the dephasing of the qubit in Eq. (20) with a time-dependent decay rate
@@
In the long-time limit $t\gg\tau_{\rm c}$, one recovers a constant decay rate $\gamma(t)\approx S(0)/4=c\tau_{\rm c}^{2}/4$, which connects to our previous discussion of the effective exponential decay of the Ramsey signal $p^{\rm TL}_{i}(m_{x})\approx\big{(}1+(-1)^{m_{x}}\mathord{\rm e}^{-t_{i}/T_{2}}\big{)}/2$ and the decoherence time $T_{2}=2/c\tau_{\rm c}^{2}$. On the other hand, for shorter time scales, we see the effects of the noise memory through a time-inhomogeneous evolution of the coherences that goes beyond a Lindbladian description. The time-dependent decay rate is always positive, such that the two measures of non-Markovianity in Eqs. (32)-(33) vanish exactly. The pure dephasing quantum dynamical map of a qubit subjected to OU frequency noise is thus Markovian albeit not Lindbladian.
@@
a. Frequentist Ramsey estimators. We can now evaluate the $\ell$QT cost function $\mathsf{C}^{\rm pd}_{\mathrm{TL}}(\bm{\theta})$ in Eq. (34) by substituting the attenuation factor in Eq. (38) in the likelihood function $p^{\rm TL}_{i}(m_{x}|\bm{\theta})=\big{(}1+(-1)^{m_{x}}\mathord{\rm e}^{-\Gamma(t_{i})}\big{)}/2$ after a certain set of evolution times $\{t_{i},i\in\mathbb{I}_{t}\}$, and the relative frequencies for the measurement outcomes $\tilde{f}_{i}(m_{x}\,|\bm{\theta}_{\star})$. For a Lindbladian dynamics, a single measuring time $t_{1}$ would suffice for the estimation [90], which can actually be solved for analytically in the present pure dephasing context, as discussed in Sec. I of the Supplementary Material. For the OU dephasing, this is no longer the case, and we actually need at least two times $t_{1},t_{2}$. In order to assess the performance of the frequentist minimization problem in Eq. (9) under shot noise, we numerically generate the relative frequencies $\tilde{f}_{1}(m_{x}\,|\bm{\theta}_{\star}),\tilde{f}_{2}(m_{x}\,|\bm{\theta}_{\star})$ at two instants of time by sampling the probability distribution with the actual OU parameters $\bm{\theta}_{\star}=(c_{\star},\tau_{\rm c{\star}})$ a number of times $N_{\rm shot}=N_{1}+N_{2}$. In the following, rather than learning $(c_{\star},\tau_{\rm c\star})$, we will focus on two noise parameters with units of time $(T_{2\star},\tau_{\rm c\star})$, where we recall that $T_{2\star}=2/c^{\phantom{2}}_{\star}\tau_{{\rm c}\star}^{2}$ is an effective decoherence time in the long-time limit. In Fig. 2, we present a contour plot of this two-time cost function $\mathsf{C}^{\rm pd}_{\mathrm{TL}}(\bm{\theta})$, which is actually convex and allows for a neat visualization of its global minimum. We also depict with a red cross the result of a gradient-descent minimization, where one can see that the estimates $\bm{\hat{\theta}}_{\rm F}=(\hat{T}_{2},\hat{\tau}_{\rm c})$ are close to the real noise parameters. The imprecision of the estimate is a result of the shot noise, which we now quantify.
Figure 2: Gradient descent for OU dephasing $\ell$QT: Contour plot of the Ramsey cost function $\mathsf{C}^{\rm pd}_{\mathrm{TL}}(\bm{\theta})$ of Eq. (34) as a function of the noise parameters $(T_{2},\tau_{\rm c})$, where we recall that $T_{2}=2/c\tau_{\rm c}^{2}$ plays the role of an effective decoherence time in the long-time limit. We choose two evolution times and, for illustration purposes, fix them at $t_{1}=\tau_{\rm c\star}$, $t_{2}=2\tau_{\rm c\star}$, fixing the OU diffusion constant such that $T_{2\star}/\tau_{\rm c\star}=1$. We distribute $N_{\rm shot}=2\times 10^{3}$ shots equally per time step. The gradient descent of this convex problem converges towards the global minimum, and is marked with a red cross at $T_{2}/T_{2\star}\approx 0.95$ and $\tau_{\rm c}/\tau_{\rm c\star}\approx 1.07$, lying close to the real noise parameters $\bm{\theta}\mapsto\bm{\theta}_{\star}$.